# Post-detection inference for sequential changepoint localization

### v4

@@ D.1 A concrete example: Gaussian mean shift problem
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. Note that $\mathcal{C}^{\prime\prime}$ in (S8) is computable if similar conditions as listed in Section C.1 are met. For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Section H.4.
D.2 Another concrete example: Poisson rate change problem
Here we assume the pre and post-change model to be $\mathcal{P}_{i}=\{\text{Pois}(\theta):\theta\in\Theta_{i}\}$, for $i=0,1$ respectively.
In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (S13) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
We fix $t\in\{1,\cdots,\tau\}$. Let $\lambda_{t}=\max\{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$. We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$. For $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise. }\end{cases}$ (S14)
Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in earlier, with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.
If $t_{2}^{j}<t$, we have $U_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified as below:
$U_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\begin{subarray}{c}\theta\in{A}^{i,j}_{t,t^{\prime}}\\ \theta^{\prime}\in{B}^{i,j}_{t,t^{\prime}}\end{subarray}}\frac{L_{i}\left(\frac{\sum_{k=1}^{i-1}Y_{k}^{j}(\theta,\theta^{\prime})}{i-1},\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-i+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}{L_{t}\left(\frac{\sum_{k=1}^{t-1}Y_{k}^{j}(\theta,\theta^{\prime})}{t-1},\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-t+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}.$
Here ${A}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},n\leq t-1,\inf\mathcal{S}_{t-1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}_{t-1}\}\cup\{\inf\mathcal{S}_{t-1}\}\cup\{\sup\mathcal{S}_{t-1}\}$ and ${B}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ are finite sets with at most $(2+\sum_{n=1}^{t^{\prime}}X_{n}^{j})$ many elements. All the maximum are taken over a finite set; hence, $U_{t}^{j}$ can be computed easily.
@@ Appendix E Experiments for confidence intervals for pre/post change parameters
E.1 Setting II
Table E.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta=0.05$. The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.85$, indicating that our method is reliable but conservative.
Table E.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table. (Setting II)
@@ E.1 Setting II
E.2 Setting III
Table E.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$. he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.80$.
Table E.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.85(Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
@@ E.2 Setting III
E.3 Comparison with [39]
Next, we compare our confidence interval for the post-change parameter with that proposed by [42] and discussed in [39]. We implement Wu’s interval as discussed in Chapter 4.3 of [39] (or Section of [42]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{cusum}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution. We use CUSUM [19] (the same change detector as in the last experiment of Section 6), with $d=10$ for both methods. For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$. We report the average size and the conditional and unconditional coverage rates of both methods across $500$ independent runs in Table E.3, with a target conditional coverage of $\alpha^{\prime}=0.85$. The results show that the existing method [39], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally). At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
Table E.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage $0.85$) with Wu [39] (Setting II)
@@ E.3 Comparison with [39]
Appendix F Additional experiments for composite pre/post change
In this section, we present experimental results for the alternative approach in which each set $\mathcal{S}_{t-1}$ and/or $\mathcal{S}_{\tau-t+1}^{\prime}$ is discretized using 50 equispaced grid points. Let, $\{\theta_{1}^{(0)},\cdots,\theta_{I}^{(0)}\}\subseteq\mathcal{S}_{t-1}$ and $\{\theta_{1}^{(1)},\cdots,\theta_{K}^{(1)}\}\subseteq\mathcal{S}_{\tau-t+1}^{\prime}$ be the grid points. Here, $I=K=50$.
Therefore, for $j\in\{1,\cdots,B\}$, we only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $i\in[I],k\in[K]$, define
$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S1)
We run $\mathcal{A}$ on $\{Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{n}$ and denote $\tau_{i,k}^{j}$ as the detection time. We compute the test statistics based on $\tau_{i,k}^{j}$ and $Y_{1}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)}),\cdots,Y_{\tau_{i,k}^{j}}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})$. Then, we consider the following confidence set:
$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (S2)
As a special case, when $\theta_{0}$ is known, then, $\theta_{0}=\theta_{1}^{(0)}=\cdots=\theta_{I}^{(0)}=\mathcal{S}_{t-1}.$
As shown in Table F.1, the confidence sets (S2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.3 and Table 6.5. However, this grid-based approach incurs significantly higher computational cost. Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
Table F.1: Composite pre/post change
T
Pre-change
Class
Post-change
Class
Conditional
coverage
Unconditional
coverage
Size Time (s)
100 $N(0,1)$ $\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.05 365.81
100 $N(0,1)$ $\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 24.62 360.48
100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 30.88 17572.26
100 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.14 17092.11
Appendix G Simpler confidence sets for detection algorithms controlling PFA
In this section, we shall discuss how one can construct simpler confidence sets when the change detection algorithm $\mathcal{A}$ controls the probability of false alarm (PFA) at level $\delta\in(0,1)$. Importantly, this method allows us to avoid Assumption 3.5, and our results hold in the fully nonparametric case with arbitrary pre- and post-change classes, $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, respectively.
We consider the test statistic $M_{t}$ introduced in (3.4), and define the following confidence set for $T$:
@@ Appendix G Simpler confidence sets for detection algorithms controlling PFA
which is a simplified version of Equation 3.9, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$. This avoids the need to compute $r_{t}^{*}$ (and hence $\tau_{1}^{*},\cdots,\tau_{N}^{*}$), making the method easier to implement and free of Assumption 3.5.
@@
From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha}]\leq\alpha$. Hence,
@@
A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting. By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.3 and 4.4, we obtain
@@
In the next section, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
Appendix H Some additional experimental results
H.1 Experiments varying $T$ and ARL
We have seen in deriving the asymptotic length bound that it increases with the ratio $T/A_{T}$. Now, we experimentally validate the fact by conducting the same experiment as in Setting I (known pre- and post-change) in Section 6 with varying $T$ thresholds $A$ of the CUSUM detector (larger threshold means large ARL). Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 are reported in Table H.1. The results show a clear pattern: as the ratio $T/A$ increases, the resulting confidence sets become wider, reflecting growing conservativeness.
Table H.1: Setting I: Known pre- and post-change. Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 with $F_{0}=N(0,1),F_{1}=N(1,1),N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (3.1) and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
$T$ A
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
50 100 0.98 0.91 17.28 2.58 9.33
1000 100 0.99 0.21 20.88 2.45 9.14
2000 100 1 0.07 27.02 2.60 8.35
500 1000 0.98 0.97 15.77 2.62 13.22
10000 1000 1 0.27 24.29 2.36 13.25
20000 1000 1 0.07 30.16 2.29 13.22
H.2 Experiments with dependent pre- and post-change data
[…]
We use the likelihood ratio-based cusum detector:
After detection, we compute the adaptive confidence sets (4.1) with the likelihood ratio-based test statistic (3.2) and point estimate as in (3.1). Note that our adaptive method (4.1) is the only one we know that is valid for dependent data. Table H.2 contains the results averaged across $500$ runs.
Table H.2: Known pre- and post-change with dependency (Setting I)
T
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 0.99 0.98 36.12 4.35 28.99
500 0.98 0.98 35.83 4.42 27.36
H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
H.3.1 Known pre- and post-change (Setting A, PFA $\leq 0.001$)
@@ H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
with thresholds at $A=1000$. Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [35], the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct both the proposed confidence sets (universal (S1) and adaptive (S2)), with $B=100$, $\alpha=0.05$. We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.3. We also report the average absolute deviation of $\hat{T}$ (defined in (3.1)) and delay (given no false detection) in the table.
Table H.3: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
@@
Universal Adaptive Universal Adaptive (Conditional) (Conditional)
100 $N(0,1)$ $N(1,1)$ 0.95 0.94 15.44 12.60 2.74 110.91
100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 19.36 17.95 4.98 94.33
500 $N(0,1)$ $N(1,1)$ 0.97 0.97 15.92 12.65 2.24 514.12
500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 18.89 16.82 3.18 417.44
H.3.2 Known pre- and unknown post-change (Setting B, PFA $\leq 0.001$)
@@
with thresholds at $A=1000$. The weights are chosen just as described in Table 6.2. Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.075$. The test statistic and the point estimate are the same as mentioned in Setting II of Section 6. We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.4. We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection) in the table.
Table H.4: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
@@
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.82 3.68 118.76
100 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 20.15 3.36 115.64
500 $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 23.15 3.18 530.83
500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 19.27 3.23 528.92
H.3.3 Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) defined as
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (S4)
with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$. This is a stopping time of an “e-process” [20] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. The weights are chosen just as described in Table 6.3. After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.1$. The test statistic and the point estimate are the same as mentioned in Setting III of Section 6. We report the average size and coverage rates across $500$ independent runs in Table H.5. We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection). The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
Table H.5: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
T
Composite
pre-change
Composite
post-change

### v6

@@ D.1 A concrete example: Gaussian mean shift problem
$t_{1}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. Note that $\mathcal{C}^{\prime\prime}$ in (S21) is computable if similar conditions as listed in Section C.1 are met. For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.
@@ Appendix E Experiments for confidence intervals for pre/post change parameters
E.1 Known pre-change but unknown post-change
Table 8 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (21)) for $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta=0.05$. The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.85$, indicating that our method is reliable but conservative.
Table 8: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table. (Setting II)
@@ E.1 Known pre-change but unknown post-change
E.2 Unknown pre- and post-change
Table 9 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), and (21)), for $\theta_{0}$ and $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$. he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.80$.
Table 9: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.85(Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
@@ E.2 Unknown pre- and post-change
E.3 Comparison with Wu [WU20063625]
Next, we compare our confidence interval for the post-change parameter with that proposed by [WU20063625] and discussed in [wu2007inference]. We implement Wu’s interval as discussed in Chapter 4.3 of [wu2007inference] (or Section 3 of [WU20063625]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{CUSUM}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution. We use CUSUM [page1954continuous] (the same change detector as in the second last experiment of Section 6), with $d=10$ for both methods. For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$. We report the average size and the conditional and unconditional coverage rates of both methods across $500$ independent runs in Table 10, with a target conditional coverage of $\alpha^{\prime}=0.85$. The results show that the existing method [WU20063625], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally). At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
Table 10: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage $0.85$) with Wu [WU20063625] (Setting II)
@@ E.3 Comparison with Wu [WU20063625]
Appendix F Simpler confidence sets for detection algorithms controlling PFA
In this section, we shall discuss how one can construct simpler confidence sets when the change detection algorithm $\mathcal{A}$ controls the probability of false alarm (PFA) at level $\delta\in(0,1)$. Importantly, this method allows us to avoid Assumption 1, and our results hold in the fully nonparametric case with arbitrary pre- and post-change classes, $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, respectively.
We consider the test statistic $M_{t}$ introduced in (9), and define the following confidence set for $T$:
@@ Appendix F Simpler confidence sets for detection algorithms controlling PFA
which is a simplified version of Equation 14, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$. This avoids the need to compute $r_{t}^{*}$ (and hence $\tau_{1}^{*},\cdots,\tau_{N}^{*}$), making the method easier to implement and free of Assumption 1.
@@
From Theorem 1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha}]\leq\alpha$. Hence,
@@
completing the proof. ∎
A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting. By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 17, 19 and 20, we obtain
@@
In the next subsection, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
F.1 Experiments with detection algorithms that control PFA
F.1.1 Known pre- and post-change (Setting A, PFA $\leq 0.001$)
@@ F.1 Experiments with detection algorithms that control PFA
with thresholds at $A=1000$. Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [ville1939etude], the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct both the proposed confidence sets (universal (S26) and adaptive (S27)), with $B=100$, $\alpha=0.1$. We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table 11. We also report the average delay (given no false detection) in the table.
Table 11: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
@@
Universal Adaptive Universal Adaptive (Conditional)
100 $N(0,1)$ $N(1,1)$ 0.95 0.94 12.65 9.86 110.91
100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 14.22 11.85 94.33
500 $N(0,1)$ $N(1,1)$ 0.97 0.97 12.83 9.76 514.12
500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 14.79 12.13 417.44
F.1.2 Known pre- and unknown post-change (Setting B, PFA $\leq 0.001$)
@@
with thresholds at $A=1000$. The weights are chosen just as described in Table 2. Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct our universal confidence sets (S26) for $T$, with $\alpha=0.1$. The test statistic is the same as mentioned in Setting II of Section 6. We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table 12. We also report the average delay (given no false detection) in the table.
Table 12: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
@@
Size
Delay
100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 16.38 118.76
100 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 13.11 115.64
500 $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 16.92 530.83
500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 14.85 528.92
F.1.3 Composite pre- and post-change (Setting C)
We generate pre-change data from $N(0,1)$ and post-change data from $N(0,1)$, having a true changepoint at $T=100$ or $500$ — these are all unknown to the algorithm. We assume access to a historic/training data $\{Y_{1},\cdots,Y_{m}\}$, $m=200$. We detect change using the following change detector, as in Sections 2 and 3 of [aue2024state]:
$\tau=\min\left\{k\in\mathbb{N}:\frac{1}{\hat{\sigma}_{m}\sqrt{m}}w(m,k)\left|\sum_{j=1}^{k}(X_{j}-\bar{Y}_{m})\right|>A\right\},$ (S32)
where $\bar{Y}_{m}=\frac{1}{m}\sum_{i=1}^{m}Y_{i}$ is the sample mean of the training data, $\hat{\sigma}_{m}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(Y_{i}-\bar{Y}_{m})^{2}}$ is the estimated standard deviation from the training data, $w(m,k)=\left(1+\frac{k}{m}\right)^{-1}$ and $A$ is the threshold, which we vary in our experiment. After detection, we construct our universal confidence sets (S26) for $T$, with $\alpha=0.1$, $\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.25\}$ and $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$. The test statistic is the same as mentioned in Setting III of Section 6. We report the average size and coverage rates across $500$ independent runs in Table 13. We also report the average delay (given no false detection).
Table 13: Composite pre- and post-change (Setting C)
T A